OpenStudy (anonymous):

Find the following limits: a)f(x)= {6-x^2 if x1 }, lim(x->1) f(x)=

OpenStudy (anonymous):

piecewise function right? this is much easier than you think replace \(x\) by 1 in both expressions if you get the same answer, that is the limit if you do not get the same answer, there is no limit

OpenStudy (anonymous):

\[f(x)= \left\{\begin{array}{rcc} 6-x^2& \text{if} & x\leq 1 \\ 2x-1& \text{if} & x >1 \end{array} \right. \] in the first one you get \(6-1=5\) in the second you get \(2-1=1\) and since \(5\neq 1\) there is no limit

OpenStudy (anonymous):

oh ok!

OpenStudy (anonymous):

I have more can you help?

OpenStudy (anonymous):

sure ask away !

OpenStudy (anonymous):

btw the last one was easy when you knew what to do, yes?

OpenStudy (anonymous):

yeah i know i just get confused on how to go about doing them

OpenStudy (anonymous):

\[\lim_{x \rightarrow 2} (\sqrt{2x+2}-\sqrt{6}) \div 2x-4\]

OpenStudy (anonymous):

it is that part in parentheses over the part behind the division sign but i didnt know how to draw that

OpenStudy (anonymous):

first try replacing \(x\) by 2 unfortunately if you do this, you get \(\frac{0}{0}\) so we will have to do more work

OpenStudy (anonymous):

\[\lim_{x\to 2}\frac{\sqrt{2x+2}-\sqrt{6}}{2x-4}\] is your question yes?

OpenStudy (anonymous):

Yeah! sorry my computer was freaking out

OpenStudy (anonymous):

ok here is what you need to do multiply top and bottom by the "conjugate" of the numerator the conjugate of \(\sqrt{2x+2}-\sqrt{6}\) is \(\sqrt{2x+2}+\sqrt{6}\) it is similar to "rationalizing the denominator" except this time you are rationalizing the numerator

OpenStudy (anonymous):

uh huh

OpenStudy (anonymous):

leave the denominator in parentheses the reason this works is that \((\sqrt{a}+\sqrt{b})(\sqrt{a}-\sqrt{b})=a-b\) you get \[\frac{\sqrt{2x+2}-\sqrt{6}}{2x-4}\times \frac{\sqrt{2x+2}+\sqrt{6}}{\sqrt{2x+2}+\sqrt{6}}\]

OpenStudy (anonymous):

\[\frac{(2x+2-6}{(2x-4)(\sqrt{2x+2}+\sqrt{6})}=\frac{2x-4}{(2x-4)(\sqrt{2x+2}+\sqrt{6})}\] cancel, then replace \(x\) by 2

OpenStudy (anonymous):

let me know what you get hope the steps are clear

OpenStudy (anonymous):

1/0

OpenStudy (anonymous):

no

OpenStudy (anonymous):

1/sqrt 12

OpenStudy (anonymous):

careful here !!

OpenStudy (anonymous):

\[\sqrt{6}+\sqrt{6}\neq \sqrt{12}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

don't do all that work and then mess up with simple combining like terms, or your teacher will think you are not ready for the class. \[\sqrt{6}+\sqrt{6}=2\sqrt{6}\] so your answer is \(\frac{1}{2\sqrt{6}}\)

OpenStudy (anonymous):

clear or no?

OpenStudy (anonymous):

oh!

OpenStudy (anonymous):

whew like \[\sqrt{25}+\sqrt{25}=5+5=10\] not \[\sqrt{25}+\sqrt{25}=\sqrt{50}\]!!!

OpenStudy (anonymous):

yeah i get it now

OpenStudy (anonymous):

good!

OpenStudy (anonymous):

is that the limit or do we have to do something to get rid of the sqrt in the denominator

OpenStudy (anonymous):

no that is the limit the trick was to get it out of looking like \(\frac{0}{0}\) by cancelling

OpenStudy (anonymous):

that is always the trick

OpenStudy (anonymous):

oh ok well the next one is : \[\lim_{x \rightarrow -2} (x ^{2}-5x+3)/ x-1=\]

OpenStudy (anonymous):

working on it

OpenStudy (anonymous):

what happens when you replace \(x\) by \(-2\) ?

OpenStudy (anonymous):

whatever you get, that is your answer

OpenStudy (anonymous):

because the denominator will be \(-2-1=-3\) and so long as the denominator is not zero, you can just evaluate

OpenStudy (anonymous):

17/-3

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

i don't know i didn't do it but the -3 is right give me a second

OpenStudy (anonymous):

yeah numerator is 17

OpenStudy (anonymous):

yay! ok i started the next one and it comes out to 0/0 and im not sure where to go from there: \[\lim_{x \rightarrow 2} ( x ^{2}-5x+4) / x-4\]

OpenStudy (anonymous):

ok gimmick is as before factor and cancel

OpenStudy (anonymous):

is the answer 3?

OpenStudy (anonymous):

and don't think too hard in factoring since if you replace \(x\) by 4 in the numerator, you know it must factor as \[x^2-5x+4=(x-4)(\text{something})\]

OpenStudy (anonymous):

factor, cancel, you get \(x-1\) replace \(x\) by 4 and get 3 yes, answer is 3

OpenStudy (anonymous):

yay! im actually getting it this is a big thing for me! i am terrible at math but i have to take this so i am trying my best to get through it

OpenStudy (anonymous):

\[\lim_{x \rightarrow -4} (3x+12)/\left| x+4 \right|\]

OpenStudy (anonymous):

forget this one

OpenStudy (anonymous):

ok i got 0/0 to begin with but with the absolute values does that mean that the -4 turns positive?

OpenStudy (anonymous):

\(|x+4|\) is a piecewise function, so the limit from the left and from the right will be different

OpenStudy (anonymous):

no it does not

OpenStudy (anonymous):

did you mean forget about this one? like as in the absolute value?

OpenStudy (anonymous):

i mean forget about it having a limit

OpenStudy (anonymous):

it will not exist here is a nice picture to show you why http://www.wolframalpha.com/input/?i=%283x%2B12%29%2F |x%2B4| notice the nice vertical line at \(x=-4\)

OpenStudy (anonymous):

you will have to cut and paste not click absolute value signs ruin it for some reason

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

ok well i still have more if you are not tired of helping me yet :)

OpenStudy (anonymous):

the reason is this \(|x+4|\) is a piecewise function if \(x<-4\) then \(|x-4|=-x-4\) and if \(x>-4\) \(|x+4|=x+4\) you are trying to take the limit as \(x\to -4\) but the function changes definition there

OpenStudy (anonymous):

go ahead and post, but put it in a new thread, this is getting too long